OpenStudy (anonymous):

find the value in simplest form:

OpenStudy (anonymous):

\[\sqrt{8} \over 3 \sqrt{4}\]

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {\sqrt{8} \over 3\sqrt{4} } }\) \(\Large \color{Black}{\Rightarrow {\sqrt{4} \times \sqrt{8} \over 12} }\) \(\Large \color{Black}{\Rightarrow {\sqrt{32} \over 12} }\) \(\Large \color{Black}{\Rightarrow {4\sqrt{2} \over 12} }\) \(\Large \color{Black}{\Rightarrow {\sqrt2 \over 3} }\)

OpenStudy (anonymous):

i got 1 more then thats all help i need :D

Parth (parthkohli):

Lol okayy